I am trying to figure out the following problem to say what is 'b':
1/a + 2/b = 3/c
My guess is that b = c minus a

Answer by askmemath(368)

1/a + 2/b = 3/c
2/b = 3/c - 1/a
2/b = (3a - c)/ca
2ca = b(3a-c)
b = 2ca/(3a-c)